1. Twenty motors were put on test under a high-temperature setting. The lifetimes in hours of the motors under these conditions are given below. Suppose we assume that the lifetime of a motor under these conditions, X, has a「(1,0) distribution. 8 11 13 13 18 22 22 43 4457 60 76 98 106 110 138 153 154 204 281 (a) Obtain a frequency distribution and a histogram or a stem-leaf plot of the data. Use the intervals [0, 25), [25,50), . the「(1,6) model is credible? . . Based on this plot, do you think that (b) Obtain the maximu ln likelihood estimate of θ and locate it on your plot. (c) Obtain the sample median of the data, which is an estimate of the median lifetime, of a motor. What parameter is it estimating (i.e., determine the median of X)? (d) Based on the mle, what is another estimate of the median of X?
Answer #1

ナhe.akt ne Xi~ ra,e) , is log n 0 0 -0 20 01,Na, ļot.nd tith OLO be, 53 and 5P 2 · 刁 MLE 白瓜.2--

d) Now here the sample standard deviation be calculated as follows,

X d=(X-mean(X)) d^2
1 -100.15 10030.02
4 -97.15 9438.123
5 -96.15 9244.823
21 -80.15 6424.023
22 -79.15 6264.723
28 -73.15 5350.923
40 -61.15 3739.323
42 -59.15 3498.723
51 -50.15 2515.023
53 -48.15 2318.423
58 -43.15 1861.923
67 -34.15 1166.223
95 -6.15 37.8225
124 22.85 522.1225
124 22.85 522.1225
160 58.85 3463.323
202 100.85 10170.72
260 158.85 25233.32
303 201.85 40743.42
363 261.85 68565.42
Total 2023 211110.6
Mean 101.15

So the sample variance be,

S^2=211110.6/19=11111.084

And the sample standard deviation be,

S = Vi îi î 1.084 105.41

So as we know for large sample, rac{ar{X}-mu}{S/sqrt{n}} , the standardized variable approximately follows standard normal distribution so the 95% confidence interval be,

r-1.64.- V +1.64. Vn

Putting the values we get,

105.41 ( 105쁨, 10115+1 64x ) 101.15-1.64 ×

62.495, 139.805)